Operation Summary for August 20, 2010 - August 23, 2010 |
The Integrated Luminosity for the period from 8/20/10 to 8/23/10 was 4.58 pb-1. Friday August 20 The day shift began with the Tevatron (TeV), Recycler, Antiproton Source (Pbar), Linac, Booster, Main Injector (MI), and Switchyard (SY) lines conducting studies and preparing for beam, and with the NuMI, MiniBooNE, Collider Detector at Fermilab (CDF), and D-Zero (D0) experiments conducting studies and preparing for beam. During the day shift, water group personnel continued their MI-62 LCW temperature investigation throughout the shift. TeV and Pbar beam studies took place. Pbar stacking started and Pbar personnel sent one transfer of antiprotons to Recycler for studies. Pbar halted stacking temporarily due to a quadrupole magnet temperature problem. During the evening shift, the TeV, Recycler, and Pelletron conducted beam studies. Operations prepared CDF and D0 collision halls for accesses after TeV personnel completed their studies. Pbar stacked all shift and transferred antiprotons to the Recycler after they had finished their studies. Mechanical technicians began working on the MI-62 LCW three-way valve. Saturday August 21 The midnight shift began with beam studies in every machine. During the midnight shift, the collision hall accesses finished. Operations put the TeV through a wet squeeze for the rest of the shift. Operations reported that a Booster RF station (BRF18) had tripped multiple times. Operators prepared the NuMI target enclosure for an access. During the day shift, Operations put the TeV into HEP shot setup at 9:38 AM and established store 7994 at 11 AM with an average initial luminosity (AIL) of 138.7E30. Mechanical techs reported at 11 AM, that they had repaired the MI-62 LCW three-way valve. Operations reported at 2:10 PM that the Pbar lithium lens tripped off on a ground fault and required a replacement. Pbar personnel began the lens repairs, which lasted throughout the shift. NuMI came out of access late on shift and operators began turning on power supplies. During the evening shift, Operations reported that MI injection kicker problems held off beam for the first four hours of the shift. Kicker personnel came in and replaced a trigger module. Two power supply failures delayed NuMI studies. Operations terminated store 7994 around 8 PM. Operations put the TeV into shot setup started at 11:42 PM. Sunday August 22 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with Pbar not stacking due to a lens replacement. During the midnight shift, Operations established store 7996 at 1:32 AM with an AIL of 175.3E30. Operations reported that Pbar personnel had completed their repair of the lithium lens by 1:34 AM. Operations ran a low intensity beam to check out the new lens. Operations began sending high intensity beam to Pbar at 3:48 AM. Operations terminated store 7996 at 6:36 AM. Operators prepared CDF, D0, and the TeV enclosures for accesses. Operators also prepared the NuMI target hall for access. During the day shift, Operations reported that power supply personnel repaired a NuMI magnet power supply. NuMI personnel began aperture scans, which lasted until the end of shift. Operations reported that a MI vertical sextupole magnet power supply had tripped on a temperature interlock; it was repaired and back online by the end of shift. CDF came out of access at 3:39 PM. During the evening shift, Operations had the TeV ramping by 5:50 PM. Bulk power supply faults interrupted the NuMI beam several times; techs came in and repaired the supply. NuMI personnel completed their target scans at 8 PM. D0 came out of access at 8:47 PM. Operations put the TeV into shot setup at 11:18 PM. Monday August 23 The midnight shift began with the TeV in shot setup, Operations monitoring the stash, with Pbar stacking, and with NuMI and taking beam. During the midnight shift, Operations established store 7999 at 1:17 AM with an AIL of 178.4E30. Operations reported that the Pbar Debuncher bend bus power supply output started wandering; power supply technicians came in and worked on the supply. The NuMI extraction kicker tripped off at 7:15 AM and would not reset; Operations called in kicker personnel. The Plans for Monday The plans for today are to continue to tune, raise intensity, and to stack, stash, and store. The Run Coordinator will allow Linac and MI studies and would like beam established to MiniBooNE by later this afternoon. |
|
More Information Comments and Suggestions |
| last modified 08/24/2010 email Fermilab |
| Security, Privacy, Legal | |